• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name W. Deconinck

    Log entry time 18:16:12 on October 22, 2009

    Entry number 296129

    This entry is a followup to: 296126


    keyword=Problems ramping up Q1L and Q3R (even after manual reset in hall)

    Even after manual resets in the hall, we failed to bring up Q1L and Q3R.
    In fact, resetting Q3R never even seems to have an effect, as there is
    no 'significant force' necessary to flip the lever on the power supply
    (although the same indicator lights are on as on the Q2R right next to
    it). Since we are unlikely to take any good data tonight, we will just
    keep the magnets off until tomorrow morning. The techs will hopefully be
    able to resolve this problem then.